Shawn is 20 and his brother is 12 how many years ago was shown three times as old as his brother
Answer by ikleyn(52788)

.
Let x = how many years ago.


Then the condition requires  (20-x) = 3*(12-x)  ====>


20 - x = 36 - 3x  ====>  3x - x = 36 - 20  ====>  2x = 16  ====>  x = 8.